# On Subjective Uncertainty Quantification and Calibration in Natural Language Generation

### v1

@@ 4 Conclusion
This work studies uncertainty quantification in free-form natural language generation. Through a Bayesian decision-theoretic perspective, we derived principled methods to quantify the model’s task-specific subjective uncertainty and to evaluate its calibration. The proposed methods are also connected to previous work, which provides additional justification for their adoption. While the discussions (in particular those in §2.2 and §2.3) are not necessarily novel outside the context of language modelling, within this domain the decision-theoretic perspective appears new and addresses important conceptual challenges in uncertainty quantification. Experiments demonstrate that the methods extract meaningful uncertainty estimates from black-box LMs and quantify their calibration.
Our discussion concerns the quantification of subjective uncertainty. In applications where i.i.d. calibration samples are available, it would be beneficial to recalibrate (Marx et al.,, 2022) the uncertainty measures, which can be achieved through (6), or to incorporate the uncertainty measures in conformal prediction (Quach et al.,, 2023). As the capabilities of LMs continue to evolve, it is also conceivable that future-generation models may find use in “one-of-its-kind” applications where post-calibration and conformal approaches are inapplicable. In such scenarios, it would be crucial to have principled measures for subjective uncertainty, as well as the tools to dissect and understand the sources of this uncertainty, e.g., through the proposed decomposition into epistemic and aleatoric uncertainty. It would also be interesting for future work to investigate the consistency between the LM’s verbalized uncertainty (Kadavath et al.,, 2022) and the subjective uncertainty defined by its predictive distributions.
Appendix A Proof for the Equivalence between (7) and (7)
@@
Appendix B Additional Experimental Details
In both experiments we use the following versions of LM APIs: gpt-3.5-turbo-0125, gemini-1.5-flash-001, gemini-1.5-pro-001.
The machine translation experiment is implemented as follows: for each test query $x_{*}$, we sample the ICL demonstrations $(x_{1:n},y_{1:n})$, a single set of $x_{n+1:N}$, and $5$ sets of $y_{n+1:N}\sim p_{M}(x_{1:N},y_{1:n})$; then for each path sample $y_{n+1:N}$ we draw $8$ samples from $p_{M}(y_{*}\mid x_{1:N},y_{1:N},x_{*})$. The action space is defined using $5\times 1$ samples from each path sample, and the expectations in (2) and (7) using the remaining $5\times 7$ samples. For each $x_{*}$, we reuse the same set of samples for $(x_{n+1:N},x_{1:n},y_{1:n})\mid x_{*}$ across all experiments. The completion inputs $x_{n+1:N}$ are generated using Gemini-1.5-Flash with the following prompt:
@@ Appendix B Additional Experimental Details
In the above, $x_{*}$ denotes the test query. For simplicity, we report the chrF scores evaluated using the final samples $y_{*}\mid x_{1:N},y_{1:N},x_{*}$, which is similar to the reinforced ICL method (Agarwal et al.,, 2024).
Figure 3: Machine translation: reliability diagrams for a subset of experiments. Error bar indicates 95% confidence interval for the expected observed utility.

### v2

@@ 5 Conclusion
This work studies uncertainty quantification in free-form natural language generation. Through a decision-theoretic perspective we derived principled methods to quantify the model’s task-specific subjective uncertainty and to evaluate its calibration. These methods are also connected to previous work, which provides additional justification for their adoption. While the discussions are not necessarily novel outside the context of language modelling, within this domain the decision-theoretic perspective appears new and addresses important conceptual challenges in uncertainty quantification. Experiments demonstrated the practical utility of the proposed methods.
Appendix A Deferred Proof
@@
Appendix B Experimental Details
Code and model samples for all experiments are available at https://github.com/meta-inf/suq-nlg.
B.1 Question Answering: Deferred Setup Details
We adopt the prompts in Lin et al., (2024) for generation and evaluation; see their Appendix B. For evaluation, their prompt instructs the LM to “rate the level of consistency” between a provided answer and a reference based on in-context demonstrations. For generation, we add a system prompt that instructs the LM to generate single-line responses to avoid interference with the evaluation prompt template in Lin et al., (2024); we find that without the system prompt, the more recent LMs we have evaluated have a greater tendency to generate multi-line responses. The generations can still contain multiple sentence, and we find them to be varied in format and content.
The NQOpen and TriviaQA datasets contain a number of outdated questions, for example about the current holder of a public office or the latest season of an ongoing TV series. We thus use an LM (gpt-4o) to identify and remove such questions. We provide the LM with a question and its (possibly outdated) reference answer, and asks it to evaluate whether the question may fall into the above categories; the LM is asked to deprioritize its internal knowledge over the reference answer during evaluation. The prompt is shown in Fig. 5. We note that the filtering process does not affect any finding reported in the main text; see e.g. Fig. 6 (a) and (b). we chose to apply this process so that the reported results more accurately reflect the LMs’ performance on these tasks.
We access all LMs through API services, from OpenAI (for GPT models) or Google Cloud. The API versions are as follows: gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, gemini-1.5-flash-001, gemini-1.5-pro-001, claude-3.5-sonnet@20240620 (versioning from Google cloud). We use the default sampling hyperparameters which always include a temperature of 1.
We implement the prompt-based uncertainty measures as follows: for the method of Kadavath et al., (2022) we adopt their prompt template and include model generations in the prompt, as recommended in their work; for the method of Tian et al., (2023), we use the two-stage variant (Verb. 2S top-1) which separates answer and confidence generation, so that the answer generation process can be comparable with the other methods; we use their prompt for confidence generation. For all test metrics we report bootstrap CIs computed using 10000 simulations.
All prompts and an illustrative subset of model generations can be found at the code repository.
I will provide you with a trivia question and its reference answer as of {year}. Please evaluate if the answer to this question is likely to have changed between {year} - 2023.
Examples of questions that might have changed answers include:
* Recent events or records
* The latest season of an ongoing TV series
* Current holders of public offices or titles
Use your best judgment in the evaluation. In scenarios where your internal knowledge could be inaccurate, rely on common sense and the provided reference answer. Please format your response in two lines, as follows:
1. A brief explanation of your reasoning (one sentence)
2. A "Yes" or "No" answer indicating whether the answer likely changed
Please adhere to the format of the example below:
Question: when did the botswana currency first come into circulation
Reference Answer: 1976
Reasoning: The date a currency first entered circulation is a historical fact that doesnt change.
Answer: No
Now please answer the following:
Question: {question}
Reference Answer: {answer}
Figure 5: Question answering: prompt template used to filter trivia datasets.
B.2 Question Answering: Additional Results
Full results in the setting of Fig. 13 are provided in Fig. 68, respectively, where we also provide results for AUARC. As expected, across LMs with a similar level of predictive performance, AUARC is generally correlated with calibration error: see the comparison between llama-3.1-70b and gemini-1.5-pro (or gpt-4o-mini) in Fig. 7. It is also interesting to note that in the setting of Fig. 1, the calibration of the gemma2-27b model (Gemma Team,, 2024) often improves after instruction tuning; inspection on TriviaQA and SciQ appears to show that the instruction-tuned LM has a similar level of variability in its generations, but the quality of generation has generally improved.
Fig. 9 compares the reliability diagrams obtained from different uncertainty measures, evaluated on llama-3.1-70b and gpt-4o. We can see that all uncertainty measures can be overconfident on inputs with low uncertainty, while the prompt-based measures can also be underconfident on high-uncertainty inputs.
(a) llama-3.1-8b
(b) gemma2-27b
(c) llama-3.1-70b
(d) llama-3.1-70b, without filtering for trivia datasets
Figure 6: Question answering: full results in the setting of Fig. 1.
Figure 7: Question answering: full results in the setting of Fig. 2. Error bar denotes 95% bootstrap CI.
Figure 8: Question answering: results for AUARC in the setting of Fig. 3.
(a) llama-3.1-70b, NQOpen
(b) gpt-4o, NQOpen
(c) llama-3.1-70b, TriviaQA
(d) gpt-4o, TriviaQA
(e) llama-3.1-70b, CoQA
(f) gpt-4o, CoQA
(g) llama-3.1-70b, SciQ
(h) gpt-4o, SciQ
(i) llama-3.1-70b, TruthfulQA
(j) gpt-4o, TruthfulQA
Figure 9: Question answering: reliability diagrams for the gpt-4o and llama-3.1-70b models. Note that the $x$-axis is always equivalent to the negated uncertainty measure.
B.3 Machine Translation: Setup Details
We use the prompt in Agarwal et al., (2024) for the translation task, and generate $x_{n+1:N}$ using gemini-1.5-flash with the prompt in Fig. 10. Details for the LMs involved are the same as App. B.1.
The deferral experiments involve two ICL predictors. The base predictor is implemented by sampling $n=4$ demonstrations uniformly without replacement. For the more expensive predictor we append $n^{\prime}-n=124$ demonstrations that are closest to the test query in an embedding space, where embeddings are computed using the text-embedding-004 model from Google Cloud. [^6] https://cloud.google.com/vertex-ai/generative-ai/docs/model-reference/text-embeddings-api In both cases we retrieve demonstrations from the dev split of the dataset and evaluate on the devtest split. We experimented with both (nontrivial) MBR and Gibbs decoding; [^7] Note that these predictors do not coincide with the two predictors in (5). Instead, (5) is only used as a surrogate to approximate the performance improvement in the deferral process. the former is implemented by drawing $16$ samples for each query, using half of them to define $\mathcal{Y}^{\prime}_{I}$ and half to approximate the expectation in (1). The main text reports the results for the MBR predictors, as we find them to perform better across all experiments. However, the same EU measure also works well for the Gibbs predictors, as shown in Fig. 12.
Following Robinson et al., (2023), we use the chrF2++ score implemented by Post, (2018). The p-values reported in Fig. 4 are calculated using 10000 replications for the random baseline.
@@ B.3 Machine Translation: Setup Details
Figure 10: Machine translation: prompt used to define $p_{x}$. $x_{*}$ denotes the test query.
B.4 Machine Translation: Additional Results
Table 2 reports the calibration error from all models. We can see that gemini-1.5-pro consistently demonstrates worse calibration.
Table 2: Machine translation: ECE for all models.
gemini-1.5-pro gpt-4o-mini gpt-4o
Tam 0.250 0.041 0.065
Kmr 0.356 0.017 0.010
(a) English $\to$ Tamil
(b) English $\to$ Kurdish
Figure 11: Machine translation: results for gemini-1.5-pro in the setting of Fig. 4.
(a) gemini-1.5-pro
(b) gpt-4o-mini
(c) gpt-4o
Figure 12: Machine translation: deferral curve for Gibbs predictors on the English $\to$ Kurdish task.